OpenStudy (anonymous):

two consecutive positive integers have a product of 6. what are the integers? separate the answer with a comma

OpenStudy (anonymous):

Consecutive Integers are like, 1 and 2, 2 and 3, or -1 and 0..

OpenStudy (anonymous):

Positive Integers are greater than 0..

OpenStudy (anonymous):

Let the first integer be \(x\), its consecutive integer will be either \(x-1\) or \(x+1\).. It is your choice to choose. I say, let other integer be \(x+1\)..

OpenStudy (anonymous):

Then product is \(6\).. So: \(x(x+1) = 6\) As you know : \( 2(2+1) = 2 \times 3 = 6\), so \(x=2\) and \(x+1\ = 3\)

OpenStudy (anonymous):

it was 2,3

OpenStudy (anonymous):

Yes it is..